# Communication-Efficient Wireless Federated Fine-Tuning for Large-Scale AI Models

### v1

@@ IV-C Online Stage
where $\gamma^{t}=2S^{2}H(r_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}+\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1-O^{t})^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}\!+\!\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the optimaility gap at time $t$ in Theorem 1.
@@
which is always positive due to the constraint $O_{\text{min}}<O^{t}\leq 1$. This confirms the convexity of the problem, allowing us to employ efficient convex optimization algorithms or well-established software tools such as CVX to find the global optimum.
For determining the bandwidth allocation strategy $b_{k}^{t}$, we leverage the property that only selected clients, i.e. those with $a_{k}^{t}=1$, receive nonzero bandwidth, i.e. $b_{k}^{t}>0$, which allows us to derive
@@
At each iteration, the server computes the optimal values of $b_{k}^{t}$ and $O^{t}$ and communicates these parameters to the selected clients. The clients then utilize the allocated bandwidth and parameters for data transmission. The overall procedure of the proposed WFLoRA algorithm is summarized in Algorithm 3.
Algorithm 3 Two Stage Federated Algorithm
1:Offline Stage
2:Obtain optimal LoRA rank $\bar{r}$ using Algorithm 2
3:Server broadcasts pre-trained model with LoRA rank $\bar{r}$ to all clients
4:Initialize virtual queue $Q^{0}\leftarrow 0$
5:Online Stage
6:for $t=0$ to $T-1$ do
7:    At the server:
8:     Observe current network conditions and $Q^{t}$
9:     Solve Problem $\mathcal{P}_{3}$ to obtain optimal $O^{t}$ and $b_{k}^{t}$
10:     Broadcast $O^{t}$, $b_{k}^{t}$, $\theta_{B}^{t}$, $\theta_{A}^{t}$ to selected clients
11:    At each selected client $k\in\mathcal{K}^{t}$:
12:     Initialize local LoRA matrices $\theta_{B,k}^{t}\leftarrow\theta_{B}^{t}$, $\theta_{A,k}^{t}\leftarrow\theta_{A}^{t}$
13:     Perform local training using loss function (III-A)
14:     Apply SOFT and transmit $\theta_{B,k}^{t}$ and $\theta_{A,k}^{t}$ to the server
15:    At the server:
16:     Aggregate received updates:
17:     $\theta_{B}^{t+1}\leftarrow\dfrac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{B,k}^{t}$
18:     $\theta_{A}^{t+1}\leftarrow\dfrac{N}{K^{t}}\sum_{k\in\mathcal{K}^{t}}p_{k}\theta_{A,k}^{t}$
19:     Update global model: $\theta_{r}^{t+1}\leftarrow\theta_{P}+\dfrac{\alpha}{r}\theta_{B}^{t+1}\theta_{A}^{t+1}$
20:     Update virtual queue: $Q^{t+1}\leftarrow\max(Q^{t}+D^{t}-D_{th},0)$
21:end for
@@ V Numerical Results
In this section, we present simulation results to demonstrate the effectiveness of the proposed SOFT and TSFA framework. All experiments were conducted using Python 3.8 on an Ubuntu server equipped with NVIDIA GeForce RTX 3090 GPUs. The total number of clients is $100$ in all experiments.
We evaluated the performance of our method on the CIFAR-100 datasets. The CIFAR-100 dataset consists of $60,000$ with $32\times 32$ color images in $100$ classes, with $600$ images per class. There are $50,000$ in training images and $10,000$ in test images. The classes are grouped into $20$ superclasses, each containing five classes, covering a wide range of objects such as animals, vehicles, and everyday items. In our experiments, we utilized transformer-based pre-trained models for fine-tuning. Specifically, we employed the ViT-Base model, which stands for the Vision Transformer Base model. ViT is a transformer-based architecture specifically designed for image recognition tasks.
Regarding the baseline schemes in the simulation, we consider the following frameworks:
  • Top-q LoRA (TLoRA)[17, 9]: Each client selects and transmits only the elements of the LoRA matrices with the largest magnitudes. Specifically, after computing the low-rank adaptation matrices, clients identify the top-$q$ elements with the highest absolute values and sparsify the rest.
  • Random LoRA (RLoRA)[9]: Each client selects and transmits only the elements of the LoRA matrices randomly without considering their magnitudes or positions within the matrices.
  • Structured LoRA (SLoRA)[11]: Each client selects and transmits only the elements of the LoRA matrices in a structured manner. Specifically, each client preserves the elements corresponding to lower rank indices and applies sparsification to those associated with higher rank indices. Here, rank indice refers to the ordering of LoRA components as defined by the model’s architecture.
  • Ideal LoRA (ILoRA)[3]: The Ideal LoRA framework represents an idealized environment where there are no communication constraints between clients and the server. In this scenario, clients transmit full LoRA matrices without any sparsification. This serves as an upper bound for performance.
V-A Effects of SOFT
@@
In this subsection, we evaluate the effectiveness of the SOFT method by comparing it with the baseline methods. Figs. 3(a) and 3(b) illustrate the test accuracy on the CIFAR-100 dataset using various sparsification methods with LoRA rank $r=4$ and $r=8$, respectively. We set the sparsification ratio as $0.5$ for all baselines and the proposed schemes. Notably, SOFT consistently maintains higher accuracy compared to the other schemes. This superior performance can be attributed to SOFT’s ability to dynamically sparsify the LoRA modules by considering the importance of individual parameters. In LoRA module-based fine-tuning, two matrices are multiplied to form a single layer. TLoRA exhibits lower performance in this context. This is because simply selecting the top-$k$ elements does not effectively capture the crucial interactions between the two matrices involved in LoRA. As a result, the essential information required for accurate model updates may be lost, leading to suboptimal performance. Similarly, SLoRA, which transmits only the elements associated with the lower rank indices throughout training, underperforms SOFT. Since SLoRA sets a fixed sparsity pattern without adapting to the varying importance of parameters during training, it cannot capture the dynamic changes in the model’s weight distribution. This rigidity results in less effective learning and lower accuracy. Furthermore, RLoRA is shown to be unsuitable for LoRA-based fine-tuning. Randomly selecting parameters to transmit fails to account for their significance in the model, which can hinder the learning process and prevent convergence. This approach leads to poor model performance.
@@ V-B Effects of Offline and Online Stage
In Section IV, we present an optimization framework that employs a two stage process: an offline stage for pre-training configuration and an online stage for dynamic parameter adjustment during training. In this subsection, we evaluate the performance of TSFA by comparing it with a one-stage variant (OSFA), which forgoes pre-optimization of the LoRA rank $r$ and instead relies solely on online adjustments with arbitrarily chosen $r$ values. The experiments were conducted using $N=100$, $K^{t}=10$, and $V=0.0001$ under a Rayleigh block fading channel characterized by a zero mean and unit variance, with constant channel gains during each user uplink transmission. The proposed SOFT method was utilized for sparsification.
Fig. 3(c) clearly demonstrates that pre-selecting an appropriate LoRA rank $r$ is essential for optimal performance. When OSFA with $r=1$, $r=2$ or $r=4$, the model suffers from insufficient capacity, despite benefiting from lower sparsification ratios compared to TSFA. In these cases, the reduced rank is inadequate to capture the model’s complexity, leading to under-parameterization and poor learning of the underlying data distribution. Conversely, when OSFA employs higher $r$ values such as $r=16$, $r=32$, or $r=64$, the increased model capacity requires a corresponding increase in the sparsification ratio to satisfy communication constraints. This, in turn, results in excessive sparsification, which limits the transmission of critical parameter updates and adversely affects learning. These findings underscore the importance of the offline stage in TSFA. By determining an appropriate LoRA rank $r$ prior to training, the two stage approach effectively balances model capacity and communication efficiency, thereby avoiding the pitfalls of both under-parameterization and over-sparsification.
V-C Effects of Covariance
@@ VI Conclusion
In this paper, we proposed a wireless federated LoRA fine-tuning framework that bridges the gap between state-of-the-art LLMs and the practical constraints of FL environments. Our analysis underscores the impact of LoRA rank and covariance on learning dynamics. We introduced SOFT for efficient parameter selection and TSFA for dynamically optimizing sparsification and bandwidth. By leveraging Lyapunov optimization, TSFA ensures long-term latency stability and resource efficiency. Experiments show that the proposed framework, enhanced with SOFT and TSFA, achieves performance on par with centralized methods while significantly reducing communication overhead, providing a scalable solution for large-scale FL over wireless networks.
@@ Appendix A Proof of Lemma 1
where (38) is by $\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}\leq\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ and (40) is satisfied by $\gamma=\frac{O_{k}}{2(1-O_{k})}$, which ensures $(1+\gamma)(1-O_{k})<1$ for $10<O_{k}<$ 1. Moreover, the last inequality is due to $2-O_{k}\leq 2$.
@@ Appendix B Proof of Theorem 1
where the second equality is due to the property $-\mathbf{a}^{\textsf{T}}\mathbf{b}=\frac{1}{2}(-\|\mathbf{a}\|_{2}^{2}-\|\mathbf{b}\|_{2}^{2}+\|\mathbf{a-b}\|_{2}^{2})$. Next, we use the following lemma:
@@
$\!+\!\eta(r_{\!\max}\!-\!r)\!H\!S^{2}\!W^{2}\!\!+\!2\eta S^{2}\!\phi rW^{2}\!\!+\!\frac{\!4\eta^{3}\!(N\!-\!K^{t}\!)}{K^{t}(N\!-\!1)}\!S^{2}\!E^{2}\!G^{2}$
$\!+\!\frac{2\eta N}{\!K^{t}(O^{t})^{4}}\!(1\!-\!O^{t})^{2}rS^{2}\!W^{4}\!+\!\frac{\!\eta^{3}\!E(E\!-\!1)\!(2E\!-\!1)}{12}\!S^{2}\!G^{2}.$ (56)
Averaging the above inequality over iteration from 00 to $T-1$ and using the property that $\theta_{o}^{0}=\theta_{r}^{0}$ since all LoRA modules are initialized with 0 at the initial stage, we can finally obtain the upper bound in Theorem 1.
@@ Appendix C Proof of Lemma 3
$\leq\!2S^{2}\underbrace{\mathbb{E}\!\left[\!\left\|\theta_{o}^{t}\!-\!\hat{\theta}_{r}^{t}\!\right\|_{F}^{2}\!\right]}_{(c1)}\!+\!2S^{2}\underbrace{\mathbb{E}\!\left[\!\left\|\hat{\theta}_{r}^{t}\!-\!\theta_{r}^{t}\right\|_{F}^{2}\!\right]}_{(c2)}$ (61)
@@
where $G^{2}=\max_{k}G_{k}^{2}$. Next, we use a simple inequality of $\|\mathbf{a}-\mathbf{b}\|_{F}^{2}\leq 2\|\mathbf{a}\|_{F}^{2}+2\|\mathbf{b}\|_{F}^{2}$, Then, $(c22)$ can be bounded as

### v2

@@ IV-C Online Stage
where $\gamma^{t}$ is the upper-bound on the optimality gap at time $t$ in Theorem 1.
@@
which is always positive due to the constraint $O_{\text{min}}<O^{t}\leq 1$. This confirms the convexity of the problem, allowing us to employ efficient convex optimization algorithms or well-established software tools such as CVX to find the global optimum. Once $O^{t}$ is determined, the overall delay $D^{t}$ for the current iteration is set. Based on (6) and $D^{t}=\max_{k\in\mathcal{K}^{t}}D^{t}_{k}$, we can now determine the bandwidth allocation strategy $b_{k}^{t}$. We leverage the property that only selected clients, i.e. those with $a_{k}^{t}=1$, receive nonzero bandwidth, i.e. $b_{k}^{t}>0$, which allows us to derive
@@
At each iteration, the server computes the optimal values of $b_{k}^{t}$ and $O^{t}$ and communicates these parameters to the selected clients. The clients then utilize the allocated bandwidth and parameters for data transmission.
@@ V Numerical Results
In this section, we present simulation results to demonstrate the effectiveness of the proposed SOFT and TSFA framework. All experiments were conducted using Python 3.8 on an Ubuntu server equipped with NVIDIA GeForce RTX 3090 GPUs. The total number of clients is 100, and 10 clients are uniformly sampled in each round in all experiments.
@@
We evaluated the performance of our method on the CIFAR-100 datasets [26]. In our experiments, we utilized transformer-based pre-trained models for fine-tuning. Specifically, we employed the pre-trained ViT-Base model, which stands for the Vision Transformer Base model. ViT is a transformer-based architecture specifically designed for image recognition tasks. LoRA layers are inserted into Blocks 0, 3, 6, and 9 of the ViT-Base model. Training was performed using the AdamW optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. The local batch size and the number of local epochs were set to 8 and 1, respectively. Regarding the baseline schemes in the simulation, we consider the following frameworks:
  • FLASC[19]: Each client selects and transmits only the elements of the LoRA matrices with the largest magnitudes. Specifically, after computing the low-rank adaptation matrices, clients identify the top-$q$ elements with the highest absolute values and sparsify the rest.
  • RLoRA[8]: Each client selects and transmits only the elements of the LoRA matrices randomly without considering their magnitudes or positions within the matrices.
  • HetLoRA[16]: Each client selects and transmits only the elements of the LoRA matrices in a structured manner. Specifically, each client preserves the elements corresponding to lower rank indices and applies sparsification to those associated with higher rank indices. Here, rank indice refers to the ordering of LoRA components as defined by the model’s architecture.
  • FedLoDrop[15]: Each client employs a structured random sparsification strategy on the LoRA matrices. Specifically, the same randomly selected LoRA rank components are retained in both matrices, so that the corresponding columns of the left LoRA matrix and rows of the right LoRA matrix are transmitted while the others are masked to zero. Unlike the original FedLoDrop, which uses Bernoulli sparsification and results in a stochastic transmission size, we enforce an exact communication budget by retaining precisely the number of rank components specified by the target sparsification ratio.
  • ILoRA[3]: The Ideal LoRA framework represents an idealized environment where there are no communication constraints between clients and the server. In this scenario, clients transmit full LoRA matrices without any sparsification. This serves as an upper bound for performance.
V-A Effects of SOFT
(a) Task loss only
(b) Proposed loss
(c) Test accuracy comparison between (a) and (b)
Figure 4: $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ of Vit-Base blocks.6.mlp.fc1 layer: (a) without and (b) with the proposed loss function and (c) comparision of their test accuracy.
@@ V-A Effects of SOFT
In this subsection, we evaluate the effectiveness of the SOFT method by comparing it with the baseline methods. Figs. 3(a) and 3(b) illustrate the test accuracy on the CIFAR-100 dataset using various sparsification methods with LoRA rank $r=4$ and $r=8$, respectively. We set the sparsification ratio as $0.5$ for all baselines and the proposed schemes to ensure a fair comparison under the same communication budget. Notably, SOFT consistently maintains higher accuracy compared to the other schemes. This superior performance can be attributed to SOFT’s ability to dynamically sparsify the LoRA modules by considering the importance of individual parameters. In LoRA module-based fine-tuning, two matrices are multiplied to form a single layer. FLASC exhibits lower performance in this context. This is because simply selecting the top-$q$ elements does not effectively capture the crucial interactions between the two matrices involved in LoRA. As a result, the essential information required for accurate model updates may be lost, leading to suboptimal performance.
Similarly, HetLoRA, which transmits only the elements associated with the lower rank indices throughout training, underperforms SOFT. Since HetLoRA sets a fixed sparsity pattern without adapting to the varying importance of parameters during training, it cannot capture the dynamic changes in the model’s weight distribution. This rigidity results in less effective learning and lower accuracy. Furthermore, RLoRA and FedLoDrop are not well suited for LoRA-based fine-tuning. RLoRA randomly selects parameters for transmission without considering their importance, which can impair the learning process and hinder convergence. Similarly, FedLoDrop randomly sparsifies LoRA rank components without accounting for their contribution to the overall low-rank update. As a result, the transmitted updates may fail to preserve the most informative components, leading to significant degradation in update quality.
@@ V-B Effects of Offline and Online Stage
In Section IV, we present an optimization framework that employs a two-stage process: an offline stage for pre-training configuration and an online stage for dynamic parameter adjustment during training. In this subsection, we evaluate the performance of TSFA by comparing it with a one-stage variant (OSFA), which forgoes pre-optimization of the LoRA rank $r$ and instead relies solely on online adjustments with arbitrarily chosen $r$ values. The experiments were conducted using $N=100$, $K^{t}=10$, $D_{th}=500s$ and $V=0.0001$ under a Rayleigh block fading channel characterized by a zero mean and unit variance, with constant channel gains during each user uplink transmission. The SOFT was utilized for sparsification.
Fig. 3(c) clearly demonstrates that pre-selecting an appropriate LoRA rank $r$ is essential for optimal performance. When OSFA with $r=1$, $r=2$ or $r=4$, the model suffers from insufficient capacity, despite benefiting from lower sparsification ratios compared to TSFA. In these cases, the reduced rank is inadequate to capture the model’s complexity, leading to under-parameterization and poor learning of the underlying data distribution. Conversely, when OSFA employs higher $r$ values such as $r=16$, $r=32$, or $r=64$, the increased model capacity requires a corresponding increase in the sparsification ratio to satisfy communication constraints. This, in turn, results in excessive sparsification, which limits the transmission of critical parameter updates and adversely affects learning. These findings underscore the importance of the offline stage in TSFA. By determining an appropriate LoRA rank $r$ prior to training, the two-stage approach effectively balances model capacity and communication efficiency, thereby avoiding the pitfalls of both under-parameterization and over-sparsification.
V-C Effects of Proposed Loss Function
Figs. 4(a) and 4(b) illustrate the effect of our proposed loss function on promoting near orthogonality within $\theta_{B,k}$ and $\theta_{A,k}$. After training with the proposed loss function as shown in Fig. 4(b), the product $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ becomes approximately diagonal, demonstrating the successful enforcement of nearly orthogonal properties. Furthermore, Fig. 4(c) compares the accuracy of CIFAR-100 between with and without the proposed orthogonal loss. Adding the orthogonal loss results in negligible performance degradation and yields accuracy nearly identical to that of the task-loss-only baseline. These results indicate that our method enables effective sparsification without requiring full matrix multiplication or SVD operations, thereby maintaining performance while significantly reducing computational overhead.
@@ V-D Effects of Covariance
Figure 7: Test accuracy of AG News dataset with rank $r=8$.
V-E Results on Other Datasets and Models
To verify that our approach generalizes beyond the vision domain, we further evaluate it on a representative NLP benchmark. Specifically, we fine-tune DistilBERT base-uncased, a 6-layer transformer distilled from BERT, on the AG News dataset [27]. We set the LoRA rank to $r=8$, the sparsification ratio to $O_{k}=0.5$, and maintain a consistent aggregation schedule to isolate the effect of our method. As shown in Fig. 7, our method consistently outperforms baseline schemes on this NLP task, demonstrating that the proposed technique is modality-agnostic and remains effective beyond image data. Furthermore, since our mechanism operates entirely at the LoRA update level, it can be seamlessly integrated into a wide range of LoRA-based PEFT frameworks, suggesting strong potential for on-device fine-tuning in natural language applications.
@@ VI Conclusion
In this paper, we developed a wireless federated LoRA fine-tuning framework for communication-efficient on-device adaptation of pre-trained models. Our analysis showed that structural LoRA rank selection and the covariance induced by separate aggregation of LoRA factors play important roles in convergence. Furthermore, we proposed SOFT for efficient parameter selection and TSFA for dynamically optimizing sparsification and bandwidth. By leveraging Lyapunov optimization, TSFA ensures long-term latency stability and resource efficiency. Experiments show that the proposed framework, enhanced with SOFT and TSFA, achieves performance on par with existing methods while significantly reducing communication overhead.
@@ Appendix A Proof of Lemma 1
where (48) is by $\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}\leq\max_{0\leq i\leq t}\|\Delta\tilde{\theta}_{k}^{i}\|_{F}^{2}$ and (50) is satisfied by $\gamma=\frac{O_{k}}{2(1-O_{k})}$, which ensures $(1+\gamma)(1-O_{k})<1$ for $0<O_{k}<1$.
@@ Appendix B Proof of Theorem 1
where the second equality is due to the property $\mathbf{a}^{\textsf{T}}\mathbf{b}=\frac{1}{2}(-\|\mathbf{a}\|_{2}^{2}-\|\mathbf{b}\|_{2}^{2}+\|\mathbf{a-b}\|_{2}^{2})$. Next, we use the following lemma:
@@
$+\eta(r_{\max}-r)HS^{2}W^{2}+2\eta S^{2}\phi rW^{2}$
$+\frac{4\eta^{3}(N-K^{t})}{K^{t}(N-1)}S^{2}E^{2}G^{2}+\frac{2\eta N}{K^{t}(O^{t})^{4}}(1-O^{t})^{2}rS^{2}W^{4}$
$+\frac{\eta^{3}E(E-1)(2E-1)}{12}S^{2}G^{2}.$ (66)
Averaging the above inequality over iteration from 0 to T1T-1 and using the property that $\theta_{o}^{0}=\theta_{r}^{0}$ since all LoRA modules are initialized with 0 at the initial stage, we can finally obtain the upper bound in Theorem 1.
@@ Appendix C Proof of Lemma 3
$\leq 2S^{2}\underbrace{\mathbb{E}\left[\left\|\theta_{o}^{t}-\hat{\theta}_{r}^{t}\right\|_{F}^{2}\right]}_{(c1)}+2S^{2}\underbrace{\mathbb{E}\left[\left\|\hat{\theta}_{r}^{t}-\theta_{r}^{t}\right\|_{F}^{2}\right]}_{(c2)}$ (71)
@@
where $G^{2}=\max_{k}G_{k}^{2}$. Next, we use a simple inequality of $\|\mathbf{a}\!-\!\mathbf{b}\|_{F}^{2}\!\leq\!2\|\mathbf{a}\|_{F}^{2}\!+\!2\|\mathbf{b}\|_{F}^{2}$, Then, $(c22)$ can be bounded as